How do I solve 3%of$36.00?
Answer by RAY100(1637)

Remember,,,,% means per centum or 1/100,,,,"of" means multiply
.
3% 0f $36,,,3/100 * 36 = .03*36 =$ 1.08
.
check, 1.08/36 =.03 = 3%